# ATLAS: Adaptive Topology-based Learning at Scale for Homophilic and Heterophilic Graphs

### v2

@@ 1 Introduction
Node classification, a fundamental problem in graph learning, involves identifying labels of nodes in a graph and has wide applications in many domains including social networks, citation networks, recommendation systems, knowledge graphs and bioinformatics  (Khemani, 2024; Wu et al., 2019b; Zhou et al., 2021). Accurate classification requires two complementary pieces of information–(i) the features at each node, and (ii) the connections between the node and its neighbors. Neural network methods such as Multi-Layer Perceptrons (MLPs) are fast but do not include information about the connections. Graph Neural Networks (GNNs) address this problem by aggregating the features between neighboring nodes, but the process is expensive, and difficult to scale to large graphs. Although the graph structure can be represented as feature vectors using different node embedding techniques (Perozzi et al., 2014; Grover and Leskovec, 2016; Tang et al., 2015), or through the use of community detection (Sun et al., 2019; Kamiński et al., 2024), the issue remains as to how many hops of neighbors should be considered and how fine-grained the communities should be. Larger hops or coarse grained community can lead to information smoothing, while smaller hops or fine grained communities can lead to information loss. Further, the hypothesis that aggregating features of neighbors can improve accuracy of node classification is only true for homophilic networks (where nodes of similar classes are connected). In heterophilic networks, where the connection between nodes need not imply similarity of class, this strategy leads to lower accuracy. Based on these observations, we posit, matching structural information (i.e. size of hops or communities) with how well it aligns with the classification is necessary for producing accurate results.
@@ 3 Methodology
Feature Augmentation. For a given resolution parameter $\gamma$, let the communities be $\mathbf{c}^{(\gamma)}\in\{1,\dots,k_{\gamma}\}$, and each node is assigned to one of the communities in $k_{\gamma}$. This assignment is represented as a one-hot encoded matrix $\mathbf{H}^{(\gamma)}$( equation 3). To reduce dimensionality, each one-hot matrix is projected into a dense embedding space using a trainable weight matrix $\mathbf{W}^{(\gamma)}$ ( equation 4). The embeddings from all resolutions are concatenated to form $\mathbf{E}$ ( equation 5), which is then further concatenated with the original features $\mathbf{X}$ to yield the augmented feature matrix $\mathbf{Z}$. The augmented feature matrix $\mathbf{Z}$ is fed to an MLP $f_{\theta}$ to produce logits; a task-dependent function $\phi$ (e.g., softmax or elementwise sigmoid), applied row-wise, converts them to probabilities $\hat{\mathbf{Y}}$ ( equation 7).
$\mathbf{X}\in\mathbb{R}^{n\times D},\quad\Gamma=\{\gamma_{1},\dots,\gamma_{T}\}$ (1)
@@
$\mathbf{Z}=[\mathbf{X}\;\|\;\mathbf{E}]\in\mathbb{R}^{n\times(D+Td_{c})}$ (6)
$\hat{\mathbf{Y}}=\phi\!\big(f_{\theta}(\mathbf{Z})\big)\in[0,1]^{n\times C}$ (7)
@@ 4.1 Q1: Accuracy Across Homophily Regimes
Table 1 reports results on the eight medium-sized benchmarks. We group these datasets into three structural-bias regimes—high, low, and negative structural bias—based on how informative their community structure is for the labels; we formalize this notion in Section 5. On graphs with negative structural bias, ATLAS improves over GCN by up to $20$ percentage points, and on high structural-bias graphs it improves over a feature-only MLP by more than $11$ percentage points. Although it does not attain the best score on every dataset, an appropriate choice of resolution parameters typically allows ATLAS to match or closely track the strongest baseline across both homophilic and heterophilic regimes. The main outlier is Roman-Empire, where accuracy appears to be largely driven by raw node features. FSGNN explicitly concatenates neighborhood features to strengthen this input signal. When we equip ATLAS with the same neighborhood-feature concatenation (ATLAS-NF), its accuracy on Roman-Empire rises to within roughly two percentage points of FSGNN, while also achieving the best performance on Tolokers.
[…]
@@
ATLAS 87.09 $\pm$ 1.62 82.19 $\pm$ 0.73 88.85 $\pm$ 0.48 42.76 $\pm$ 3.47 53.15 $\pm$ 0.61 38.48 $\pm$ 0.93 40.82 $\pm$ 2.12 66.22 $\pm$ 0.53
ATLAS-NF 86.73 $\pm$ 1.04 83.02 $\pm$ 0.74 88.76 $\pm$ 0.36 40.02 $\pm$ 2.79 52.30 $\pm$ 0.64 34.26 $\pm$ 0.97 36.98 $\pm$ 2.37 78.10 $\pm$ 0.59
ATLAS–MLP (pp) +11.65 +9.22 +1.60 +6.76 +13.32 +3.52 +6.53 +0.64
ATLAS–GCN (pp) +0.08 +7.26 +2.14 +5.65 +10.37 +9.99 +8.12 +20.54
ATLAS–Average (pp) +0.92 +5.97 +0.40 +4.99 +8.51 +2.13 +4.54 +1.67
Table 2: Large-graph performance. Baselines from Zeng et al. (2020); Hu et al. (2020). Bottom rows report ATLAS improvements over MLP and over GCN (absolute units). Cells highlighted in yellow indicate the best score for each dataset.
@@
MLP 0.7435 $\pm$ 0.0016 0.6106 $\pm$ 0.0008 0.4717 $\pm$ 0.0011 0.6546 $\pm$ 0.0011 0.8204 $\pm$ 0.0002
GCN 0.9330 $\pm$ 0.0001 0.7564 $\pm$ 0.0021 0.4920 $\pm$ 0.0030 0.3780 $\pm$ 0.0010 0.2810 $\pm$ 0.0050
GraphSAGE 0.9530 $\pm$ 0.0010 0.8061 $\pm$ 0.0016 0.5010 $\pm$ 0.0130 0.6340 $\pm$ 0.0060 0.7580 $\pm$ 0.0020
ClusterGCN 0.9540 $\pm$ 0.0010 0.7862 $\pm$ 0.0061 0.4810 $\pm$ 0.0050 0.6090 $\pm$ 0.0050 0.7590 $\pm$ 0.0080
GraphSAINT 0.9660 $\pm$ 0.0010 0.7536 $\pm$ 0.0034 0.5110 $\pm$ 0.0010 0.6530 $\pm$ 0.0030 0.8150 $\pm$ 0.0010
SGC 0.9351 $\pm$ 0.0004 0.6748 $\pm$ 0.0011 0.5035 $\pm$ 0.0005 0.2356 $\pm$ 0.0002 0.2262 $\pm$ 0.0028
SIGN 0.9595 $\pm$ 0.0002 0.8052 $\pm$ 0.0016 0.5160 $\pm$ 0.0011 0.5798 $\pm$ 0.0012 0.7424 $\pm$ 0.0002
SAGN 0.9648 $\pm$ 0.0003 0.8121 $\pm$ 0.0007 0.5007 $\pm$ 0.0011 0.6155 $\pm$ 0.0040 0.7682 $\pm$ 0.0115
GAMLP 0.9673 $\pm$ 0.0003 0.8376 $\pm$ 0.0019 0.5258 $\pm$ 0.0012 0.5784 $\pm$ 0.0154 0.7599 $\pm$ 0.0026
ATLAS 0.9574 $\pm$ 0.0004 0.7998 $\pm$ 0.0014 0.5156 $\pm$ 0.0014 0.6546 $\pm$ 0.0011 0.8204 $\pm$ 0.0002
ATLAS-NF 0.9410 $\pm$ 0.0007 0.7507 $\pm$ 0.0030 0.5272 $\pm$ 0.0011 0.5740 $\pm$ 0.0021 0.7783 $\pm$ 0.0010
ATLAS–MLP +0.2139 +0.1892 +0.0439 +0.0000 +0.0000
ATLAS–GCN +0.0244 +0.0434 +0.0236 +0.2766 +0.5394
ATLAS–Average +0.0267 +0.0395 +0.0153 +0.1059 +0.1615
@@ 4.2 Q2: Efficiency and Scalability on Large Graphs
Accuracy. Table 2 shows that ATLAS scales to million-node graphs and is competitive across all structural-bias regimes while consistently improving over MLP and GCN. On high structural-bias graphs, it stays close to the best methods with gains up to about +0.21 over MLP and +0.03 over GCN. On Flickr, ATLAS-NF (ATLAS with neighbor features) surpasses ATLAS, indicating the impact of enhanced feature signal from neighbours, and on negative structural-bias graphs ATLAS maintains strong performance while aggregation-heavy GNNs degrade.
@@ 5 Accuracy Under Community Refinement
  • Low structural bias (e.g., Amazon-Ratings, Chameleon-Filtered, Flickr): Community structure is only weakly label-aligned. Coarse communities capture most of this limited structural signal, and adding finer communities yields at most small additional gains. As $Q_{\min}$ decreases and more refined communities are included, ATLAS improves moderately over the MLP, while GCN shows at best small or sometimes no gains over MLP; see Figure 5 (top center and bottom center). Here, topology provides some extra information, but node features remain the primary driver of performance.

### v3

@@ 1 Introduction
Node classification, a fundamental problem in graph learning, involves identifying labels of nodes in a graph and has wide applications in many domains including social networks, citation networks, recommendation systems, knowledge graphs and bioinformatics (Khemani, 2024; Wu et al., 2019b; Zhou et al., 2021). Accurate classification requires two complementary pieces of information–(i) the features at each node, and (ii) the connections between the node and its neighbors. Neural network methods such as Multi-Layer Perceptrons (MLPs) are fast but do not include information about the connections. Graph Neural Networks (GNNs) address this problem by aggregating the features between neighboring nodes, but the process is expensive and difficult to scale to large graphs. To improve scalability, sampling-based methods construct mini-batches by sampling neighborhoods or subgraphs (e.g., GraphSAGE, ClusterGCN, GraphSAINT) (Hamilton et al., 2017; Chiang et al., 2019; Zeng et al., 2020); however, this introduces an additional granularity choice—the fan-out or subgraph/cluster size—that trades off computational cost against how much structural context is captured per update. Although the graph structure can be represented as feature vectors using different node embedding techniques (Perozzi et al., 2014; Grover and Leskovec, 2016; Tang et al., 2015), or through the use of community-aware representations/features derived from community detection (Sun et al., 2019; Kamiński et al., 2024), the issue remains as to how many hops of neighbors should be considered and how fine-grained the communities (or sampled subgraphs) should be. Larger hops, larger sampled subgraphs, or coarse-grained communities can lead to information smoothing, while smaller hops, smaller sampled subgraphs, or fine-grained communities can lead to information loss. Further, the hypothesis that aggregating features of neighbors can improve accuracy of node classification is only true for homophilic networks (where nodes of similar classes are connected). In heterophilic networks, where the connection between nodes need not imply similarity of class, this strategy leads to lower accuracy. Based on these observations, we posit that matching structural information (i.e., size of hops, sampled subgraphs/clusters, or communities) with how well it aligns with the classification task is necessary for producing accurate results.
@@ 3 Methodology
Feature Augmentation. For a resolution $\gamma\in\Gamma$, community detection on $G$ yields node assignments $\mathbf{c}^{(\gamma)}\in\{1,\dots,k_{\gamma}\}^{n}$. We one-hot encode them as $\mathbf{H}^{(\gamma)}$ (equation 2) and project to a dense space using a trainable $\mathbf{W}^{(\gamma)}$ (equation 3). Concatenating embeddings across resolutions forms $\mathbf{E}$ (equation 4), which is concatenated with node features $\mathbf{X}$ to obtain $\mathbf{Z}_{\text{ATLAS}}$. An MLP $f_{\theta}$ outputs logits, and a task-dependent map $\phi$ (e.g., softmax or elementwise sigmoid) produces probabilities $\hat{\mathbf{Y}}$ (equation 6).
@@
$\mathbf{Z}_{\text{ATLAS}}$ $=[\,\mathbf{X}\ \|\ \mathbf{E}\,]\in\mathbb{R}^{n\times(D+Td_{c})}$ (5)
$\hat{\mathbf{Y}}$ $=\phi\!\big(f_{\theta}(\mathbf{Z}_{\text{ATLAS}})\big)\in[0,1]^{n\times C}.$ (6)
Local neighbor effects.
Multi-resolution community embeddings capture informative meso-scale structure, but may under-represent strictly local signals driven by immediate neighbors (e.g., mixed labels within communities or short-range dependencies), as well as fine-grained attribute context. Without changing the ATLAS backbone, we therefore evaluate two optional neighborhood-based augmentations computed from the original adjacency $\mathbf{A}$. ATLAS-NF aggregates 1-hop neighbor features, $\mathbf{X}_{\mathrm{NF}}=\mathbf{A}\mathbf{X}$. ATLAS-LPF seeds training labels only (one-/multi-hot) and diffuses them using degree-normalized propagation $\tilde{\mathbf{A}}=\mathrm{RowNorm}(\mathbf{A})$ to obtain a soft neighborhood label prior; we concatenate NF/LPF with $\mathbf{Z}_{\text{ATLAS}}$.
ATLAS-NF: $\mathbf{X}_{\mathrm{NF}}=\mathbf{A}\mathbf{X},\qquad\mathbf{Z}_{\mathrm{ATLAS\text{-}NF}}=[\,\mathbf{Z}_{\text{ATLAS}}\|\mathbf{X}_{\mathrm{NF}}\,].$ (7)
ATLAS-LPF: $\mathbf{Y}^{(0)}=(\mathbf{m}_{\mathrm{tr}}\mathbf{1}_{C}^{\top})\odot\mathbf{Y},\qquad\mathbf{Y}^{(h)}=\tilde{\mathbf{A}}\mathbf{Y}^{(h-1)},\ h=1,\dots,H,$
$\mathbf{Y}_{\mathrm{LPF}}=\big\|_{h=1}^{H}\mathbf{Y}^{(h)}\in\mathbb{R}^{n\times(HC)},\qquad\mathbf{Z}_{\mathrm{ATLAS\text{-}LPF}}=[\,\mathbf{Z}_{\text{ATLAS}}\|\mathbf{Y}_{\mathrm{LPF}}\,].$
$\mathbf{Z}_{\mathrm{ATLAS\text{-}NF\text{-}LPF}}=[\,\mathbf{Z}_{\text{ATLAS}}\|\mathbf{X}_{\mathrm{NF}}\|\mathbf{Y}_{\mathrm{LPF}}\,].$
@@ 4.1 Q1: Accuracy Across Homophily Regimes
Table 1 summarizes results across three homophily regimes (high, low, negative). On high-homophily graphs, ATLAS improves over a feature-only MLP (e.g., +11.65 on Cora) and remains competitive with strong GNN baselines; ATLAS-NF is best on Tolokers (83.02). On low-homophily graphs, ATLAS achieves the top result on Chameleon-Filtered (42.76), while ATLAS-LPF is best on Amazon-Ratings (53.18), indicating that a local label prior can complement meso-scale community signals. On negative-homophily graphs, ATLAS consistently outperforms GCN (e.g., +20.54 on Roman-Empire) and local augmentations further help on feature-driven cases: ATLAS-NF reaches 78.10 and ATLAS-LPF-NF attains the best overall accuracy on Roman-Empire (80.80). Overall, ATLAS is robust across regimes, with NF/LPF improving cases where strictly local attribute or neighborhood-label information is critical.
@@
ATLAS 88.37 $\pm$ 0.63 82.19 $\pm$ 0.73 88.85 $\pm$ 0.48 42.76 $\pm$ 3.47 53.15 $\pm$ 0.61 38.48 $\pm$ 0.93 40.82 $\pm$ 2.12 66.22 $\pm$ 0.53
ATLAS-NF 89.20 $\pm$ 0.45 83.02 $\pm$ 0.74 88.76 $\pm$ 0.36 40.02 $\pm$ 2.79 52.30 $\pm$ 0.64 34.26 $\pm$ 0.97 36.98 $\pm$ 2.37 78.10 $\pm$ 0.59
ATLAS-LPF 88.25 $\pm$ 0.46 82.06 $\pm$ 0.61 88.80 $\pm$ 0.12 41.54 $\pm$ 3.44 53.18 $\pm$ 0.30 38.93 $\pm$ 0.65 40.08 $\pm$ 2.19 75.41 $\pm$ 2.20
ATLAS-LPF-NF 89.06 $\pm$ 0.39 82.79 $\pm$ 0.78 88.58 $\pm$ 0.29 40.48 $\pm$ 3.53 52.28 $\pm$ 0.25 36.26 $\pm$ 0.81 38.20 $\pm$ 2.35 80.80 $\pm$ 0.84
ATLAS–MLP +12.93 +9.22 +1.60 +6.76 +13.32 +3.52 +6.53 +0.64
ATLAS–GCN +1.36 +7.26 +2.14 +5.65 +10.37 +9.99 +8.12 +20.54
ATLAS–Average +2.20 +5.97 +0.40 +4.99 +8.51 +2.13 +4.54 +1.67
Table 2: Large-graph performance. Baselines from Zeng et al. (2020); Hu et al. (2020). Bottom rows report ATLAS improvements over MLP and over GCN (absolute percentage points). Cells highlighted in yellow indicate the best score for each dataset.
@@
MLP 74.35 $\pm$ 0.16 61.06 $\pm$ 0.08 47.17 $\pm$ 0.11 65.46 $\pm$ 0.11 82.04 $\pm$ 0.02
GCN 93.30 $\pm$ 0.01 75.64 $\pm$ 0.21 49.20 $\pm$ 0.30 37.80 $\pm$ 0.10 28.10 $\pm$ 0.50
GraphSAGE 95.30 $\pm$ 0.10 80.61 $\pm$ 0.16 50.10 $\pm$ 1.30 63.40 $\pm$ 0.60 75.80 $\pm$ 0.20
ClusterGCN 95.40 $\pm$ 0.10 78.62 $\pm$ 0.61 48.10 $\pm$ 0.50 60.90 $\pm$ 0.50 75.90 $\pm$ 0.80
GraphSAINT 96.60 $\pm$ 0.10 75.36 $\pm$ 0.34 51.10 $\pm$ 0.10 65.30 $\pm$ 0.30 81.50 $\pm$ 0.10
SGC 93.51 $\pm$ 0.04 67.48 $\pm$ 0.11 50.35 $\pm$ 0.05 23.56 $\pm$ 0.02 22.62 $\pm$ 0.28
SIGN 95.95 $\pm$ 0.02 80.52 $\pm$ 0.16 51.60 $\pm$ 0.11 57.98 $\pm$ 0.12 74.24 $\pm$ 0.02
SAGN 96.48 $\pm$ 0.03 81.21 $\pm$ 0.07 50.07 $\pm$ 0.11 61.55 $\pm$ 0.40 76.82 $\pm$ 1.15
GAMLP 96.73 $\pm$ 0.03 83.76 $\pm$ 0.19 52.58 $\pm$ 0.12 57.84 $\pm$ 1.54 75.99 $\pm$ 0.26
ATLAS 95.74 $\pm$ 0.04 79.98 $\pm$ 0.14 51.56 $\pm$ 0.14 65.46 $\pm$ 0.11 82.04 $\pm$ 0.02
ATLAS-NF 94.10 $\pm$ 0.07 75.07 $\pm$ 0.30 52.72 $\pm$ 0.11 57.40 $\pm$ 0.21 77.83 $\pm$ 0.10
ATLAS-LPF 96.74 $\pm$ 0.03 80.52 $\pm$ 0.13 51.92 $\pm$ 0.14 60.20 $\pm$ 1.34 56.64 $\pm$ 0.28
ATLAS-LPF-NF 94.47 $\pm$ 0.10 75.14 $\pm$ 0.30 52.82 $\pm$ 0.16 51.92 $\pm$ 2.07 53.70 $\pm$ 0.48
ATLAS–MLP +21.39 +18.92 +4.39 +0.00 +0.00
ATLAS–GCN +2.44 +4.34 +2.36 +27.66 +53.94
ATLAS–Average +2.67 +3.95 +1.53 +10.59 +16.15
@@ 4.2 Q2: Efficiency and Scalability on Large Graphs
Accuracy.
Table 2 shows that ATLAS remains competitive at million-node scale while maintaining our propagation-free inference design: prediction is performed by an MLP over topology-derived features (multi-resolution community summaries), optionally augmented with precomputed local signals. On high structural-bias graphs, ATLAS attains $95.74$ (Reddit) and $79.98$ (ogbn-products), and ATLAS-LPF further improves performance to $96.74$ and $80.52$ by injecting a degree-normalized local label prior from training labels. On Flickr (low structural bias), local attribute context is more beneficial: ATLAS-NF improves over ATLAS ($52.72$ vs. $51.56$) and ATLAS-LPF-NF performs best among ATLAS variants ($52.82$). In contrast, on Yelp (negative structural bias), neighborhood aggregation is less label-aligned and NF/LPF reduce accuracy (ATLAS: $65.46$ vs. $57.40$ for ATLAS-NF and $60.20$ for ATLAS-LPF). Overall, these results support the structural-bias perspective: community features provide a robust meso-scale signal, while NF/LPF offer targeted gains when local neighborhood information is label-informative, without changing the ATLAS inference pipeline.
@@ 5 Accuracy Under Community Refinement
  • Low structural bias (e.g., Amazon-Ratings, Chameleon-Filtered, Flickr): Community structure is only weakly label-aligned. Coarse communities capture most of this limited structural signal, and adding finer communities yields at most small additional gains. As $Q_{\min}$ decreases and more refined communities are included, ATLAS improves moderately over the MLP, while GCN shows at best small or sometimes no gains over MLP; see Figure 5 (top center and bottom center).